OpenStudy (anonymous):

in ∫ (x^(1/3)) / ( (x^(1/3)) -1), what should i substitute?

OpenStudy (anonymous):

Quotient rule?

OpenStudy (dumbcow):

u = x^(1/3) - 1 --> x = (u+1)^3 du = dx/(3x^(2/3))

OpenStudy (dumbcow):

you could also do it another way u= x^1/3 --> x=u^3 du = dx/(3x^(2/3)) yielding 3*integral u^3/(u-1) du